Lake Power Sports sells jet skis and other powered recreational equipment. Customers pay one-third of the sales price of a jet ski when they initially purchase the ski, and then pay another one-third each year for the next two years. Because Lake has little information about the ability to collect these receivables, it uses the installment sales method for revenue recognition. In 2020, Lake began operations and sold jet skis with a total price of $630,000 that cost Lake $315,000. Lake collected $210,000 in 2020, $210,000 in 2021, and $210,000 in 2022 associated with those sales. In 2021, Lake sold jet skis with a total price of $2,070,000 that cost Lake $1,242,000. Lake collected $690,000 in 2021, $610,000 in 2022, and $610,000 in 2023 associated with those sales. In 2023, Lake also repossessed $160,000 of jet skis that were sold in 2021. Those jet skis had a fair value of $60,000 at the time they were repossessed.

In its December 31, 2021, balance sheet, Lake would report:

Multiple Choice

  • Installment receivables (net) of $933,000.

  • Deferred gross profit of $860,000.

  • Installment receivables (net) of $630,000.

  • Deferred gross profit of $1,755,000.
Answer #1

For year 2020 =

Profit rate is 50%

Collected amount = 210000 +210000 = 420000

Uncollected amount = 210000

Uncollected amount without profit = 210000 *0.50 = 105000

For year 2021 =

Profit rate is 40%

Uncollected amount = 2070000 - 690000 = 1,380,000

Uncollected amount without profit = 1,380,000 *0.60 = 828000

Installment receivables (net) of = 828000 +105000 = $933,000.